OpenStudy (anonymous):

Ill be honest i have no idea abou this one Find the product of (x + 3)(x − 3). x2 − 6x + 9 x2 + 6x + 9 x2 + 9 x2 − 9

OpenStudy (kinggeorge):

\[(x+3)(x-3)=x(x-3)+3(x-3)\]Can you finish it from there?

OpenStudy (anonymous):

x2 + 6x + 9 this one?

OpenStudy (kinggeorge):

Not quite.\[x(x−3)+3(x−3)=x^2-3x+3x-9=x^2-9\]

OpenStudy (anonymous):

16x2 − 9y2 i am pretty sure its this one maybe......

OpenStudy (kinggeorge):

You want to factor \(16x^2-9y^2\) correct?

OpenStudy (anonymous):

x2 − 6x + 9 ok i am really sure this is it :p lol

OpenStudy (kinggeorge):

First notice that both terms are perfect squares. Then, by the pattern we can observe from solving \((x+3)(x-3)=x^2-9\). Thus, \[16x^2-9y^2=(4x+3y)(4x-3y)\]

OpenStudy (kinggeorge):

And \((x+3)(x-3)=x^2-9\)